# Predicate Invention from Pixels via Pretrained Vision-Language Models

### v1

@@ A.5 Additional Prompting Details
Here, we provide additional details how we prompt a VLM for both labeling and proposal of atoms.
Atom Labeling.
Recall from Section 4.1 that the goal of atom labeling is to obtain the truth value of a particular ground atom $\underline{\psi}(o_{0},\ldots,o_{l})$ given an image-based state at timestep $t$ in a trajectory $x_{t}^{\text{img}}$. We do this by prompting a VLM with a text prompt, as well as a string representation of the atom (e.g. Cooked(patty1)) and asking it to output the truth value of the atom. Since we usually want to query for the values of many atoms at once, we provide all atoms in one single query and ask a VLM to label their truth values simultaneously.
We use a different prompt depending on whether timestep $t=0$ or $t>0$. For $t=0$, we simply provide the initial image-based state $x_{0}^{\text{img}}$ and a prompt (shown below) that asks the model to label all the values of atoms listed in the prompt with either “True’, “False” or “Unknown”.
You are a vision system for a robot. Your job is to output the values of the following predicates based on the provided visual scene. For each predicate, output True, False, or Unknown if the relevant objects are not in the scene or the value of the predicate simply cannot be determined. Output each predicate value as a bulleted list with each predicate and value on a different line. For each output value, provide an explanation as to why you labelled this predicate as having this particular value. Use the format: $<$predicate$>$: $<$truth_value$>$. $<$explanation$>$.
Predicates:
We then specify the ground atom strings whose values we’d like to query. For instance, in the “Bigger Burger” task, these might be:
@@
We found two aspects of this were critical to labeling accuracy. Firstly, chain-of-thought prompting (Wei et al. 2022) was extremely useful: our prompt asks the VLM to explicitly provide reasoning for its choices in addition to labeling the truth value. Secondly, recall that the image-based state $x^{\text{img}}$ includes segmentation of object names rendered onto the image itself (as depicted in Figure 2). This is a form of set-of-marks prompting (Yang et al. 2023) and is critical to the VLM being able to identify objects being referred to in the text prompt.
@@
For all states $x_{t}$ in a trajectory that are not the initial state (i.e., $t>0$), we prompt the VLM with the following information: (1) the image-based state from the current and previous timesteps (i.e. $x_{t}^{\text{img}}$ and $x_{t-1}^{\text{img}}$), (2) the action executed to get to the current state (i.e. $a_{t-1})$, and (3) the response of the VLM from the previous timestep.
We modify the prompt to the VLM to be as follows (here we show some of the predicates queried and a particular action for a certain timestep in a trajectory for a Burger task):
@@
You are a vision system for a robot. You are provided with two images corresponding to the states before and after a particular skill is executed. You are given a list of predicates below, and you are given the values of these predicates in the image before the skill is executed. Your job is to output the values of the following predicates in the image after the skill is executed. Pay careful attention to the visual changes between the two images to figure out which predicates change and which predicates do not change. For the predicates that change, list these separately at the end of your response. Note that in some scenes, there might be no changes. First, output a description of what changes you expect to happen based on the skill that was just run, explicitly noting the skill that was run. Second, output a description of what visual changes you see happen between the before and after images, looking specifically at the objects involved in the skill’s arguments, noting what objects these are. Next, output each predicate value in the after image as a bulleted list with each predicate and value on a different line. For each predicate value, provide an explanation as to why you labeled this predicate as having this particular value. Use the format: $<$predicate$>$: $<$truth_value$>$. $<$explanation$>$.
Your response should have three sections. Here is an outline of what your response should look like:
[START OUTLINE]
# Expected changes based on the executed skill
[insert your analysis on the expected changes you will see based on the skill that was executed]
# Visual changes observed between the images
[insert your analysis on the visual changes observed between the images]
# Predicate values in the after image
[insert your bulleted list of ‘* ¡predicate¿: ¡truth value¿. ¡explanation¿‘]
[END OUTLINE]
Predicates: available(robot)
busy(robot)
cooked(patty1)
empty(bottom_bun1)
empty(grill)
empty(robot)
empty_grill(grill)
free(robot)
Skill executed between states: Place[robot:robot, top_bun1:top_bun, patty1:patty]
We also provide the response of the VLM from a previous timestep, an example of which is shown below.
Predicate values in the first scene, before the skill was executed: # Expected changes based on the executed skill
The skill executed is ”Pick[robot:robot, top_bun1:top_bun]”. We expect the robot to pick up the top bun. The top bun should no longer be on the ground or table and should be in the robot’s possession.
# Visual changes observed between the images
In the before image, the top bun is on the ground. In the after image, the robot is holding the top bun, indicating that the robot has successfully picked it up.
# Predicate values in the after image
* available(robot): Unknown. The previous value was unknown, and the pick action does not provide information about availability.
* busy(robot): Unknown. The previous value was unknown, and the pick action does not provide information about busyness.
* cooked(patty1): True. The patty remains cooked, as it was cooked before the action. Previously, the value was true.
* empty(bottom_bun1): False. The bottom bun remains unchanged. Previously, the value was false.
* empty(grill): True. The grill remains empty, as the action does not involve the grill. Previously, the value was true.
* empty(robot): False. The robot is now holding the top bun. Previously, the value was true.
* empty_grill(grill): True. The grill remains empty, as the action does not involve the grill. Previously, the value was true.
* free(robot): Unknown. The previous value was unknown, and the pick action does not provide information about freedom. …
We found that prompting the VLM to describe differences between two scenes in a chain-of-thought fashion led to much more accurate output than simply labeling atoms based on a single scene.
@@
There are two additional prompting techniques we found helpful and important for labeling accuracy: (1) asking the VLM to “double-check” its label output based on the explanations it provided, and (2) augmenting the image-based state to include a close-up crop of the image for the objects the robot is currently interacting with. For this double-checking, we provide the VLM with the following prompt:
@@
We use double-checking only at training time to maximize labeling accuracy before hill-climbing. We augment images with additional object crops only for tasks in the Burger domain.
Atom Proposal
@@
Once again, we found that chain-of-thought prompting was critical to good proposals. We also found that ensuring all the image-based states have segmented objects with names rendered on them was important to get a good set of proposals.
@@
### Predicates for Each Action
1. **Pick[robot:robot, patty1:patty]**
- Before: on(patty1, ground)
- After: holding(robot, patty1)
- Synonyms: grasp(robot, patty1)
- Antonyms: release(robot, patty1)
2. **Place[robot:robot, patty1:patty, grill:grill]**
- Before: holding(robot, patty1)
- After: on(patty1, grill)
- Synonyms: position(patty1, grill)
- Antonyms: off(patty1, grill)
3. **Cook[robot:robot, patty1:patty, grill:grill]**
- Before: raw(patty1)
- After: cooked(patty1)
- Synonyms: grill(patty1)
- Antonyms: uncooked(patty1)
4. **Pick[robot:robot, patty1:patty]**
- Before: on(patty1, grill)
- After: holding(robot, patty1)
- Synonyms: grasp(robot, patty1)
- Antonyms: release(robot, patty1)
5. **Place[robot:robot, patty1:patty, bottom_bun1:bottom_bun]**
- Before: holding(robot, patty1)
- After: on(patty1, bottom_bun1)
- Synonyms: position(patty1, bottom_bun1)
- Antonyms: off(patty1, bottom_bun1)
6. **Pick[robot:robot, top_bun1:top_bun]**
- Before: on(top_bun1, ground)
- After: holding(robot, top_bun1)
- Synonyms: grasp(robot, top_bun1)
- Antonyms: release(robot, top_bun1)
7. **Place[robot:robot, top_bun1:top_bun, patty1:patty]**
- Before: holding(robot, top_bun1)
- After: on(top_bun1, patty1)
- Synonyms: position(top_bun1, patty1)
- Antonyms: off(top_bun1, patty1)
### Other Important Predicates
- **Initial State:**
- on(patty1, ground)
- on(bottom_bun1, ground)
- on(top_bun1, ground)
- empty(robot)
- **Final State:**
- on(top_bun1, patty1)
- on(patty1, bottom_bun1)
- empty(robot)
- **Synonyms:**
- empty(robot), free(robot)
- on(obj1, obj2). above(obj1, obj2)
- **Antonyms:** - empty(robot), holding(robot, obj)
- on(obj1, obj2), off(obj1, obj2)
@@ A.6 Additional Experimental Details
Additional Environment Details.
@@
        KettleBoiling(?k: kettle, ?kn: knob, ?b burner): is true only if ?kn is turned on, and the kettle ?k is ontop of ?b, and if ?kn and ?b are ’linked’ (i.e., ?kn is the knob that causes burner ?b to glow red with heat).
@@
      Cook(?r: robot, ?p: patty, ?g: grill). Given the robot is adjacent to grill ?g and patty ?p is atop the grill, cooks the patty so it appears grilled.
@@
      Chop(?r: robot, ?l: lettuce, ?c: cutting_board). Given the robot is adjacent to cutting board ?c and lettuce ?l is atop the cutting board, chops the lettuce with a knife that’s on the cutting board so it appears chopped.
@@
          SomewhereAboveAndPrepped(?l: lettuce, ?b: bottom_bun): turns true when ?l is somewhere above ?b and ?l is choppeed.

### v4

@@ A.5 Additional Prompting Details
Here, we provide additional details about how we prompt a VLM for both labeling and proposal of atoms.
A.5.1 Atom Labeling.
Recall from Section 4.2 that the goal of atom labeling is to obtain the truth value of a particular ground atom $\underline{\psi}(o_{0},\ldots,o_{l})$ given an image-based state at timestep $t$ in a trajectory $s_{t}^{\text{img}}$. We do this by prompting a VLM with a text prompt that includes a string representation of the atom (e.g. Cooked(patty1)), along with the relevant images in $s_{t}^{\text{img}}$, and asking it to output the truth value of the atom. Since we usually want to query for the values of many atoms at once, we provide all atoms in one single query and ask a VLM to label their truth values in batch. Having separate queries to label each ground atom would probably lead to more accurate labels, but would be much slower.
We use a different prompt depending on whether timestep $t=0$ or $t>0$. For $t=0$, we simply provide the initial image-based state $s_{0}^{\text{img}}$ and a prompt (shown below) that asks the model to label all the values of atoms listed in the prompt with either “True’, “False” or “Unknown”.
You are a vision system for a robot. Your job is to output the values of the following predicates based on the provided visual scene. For each predicate, output True, False, or Unknown if the relevant objects are not in the scene or the value of the predicate simply cannot be determined. Output each predicate value as a bulleted list with each predicate and value on a different line. For each output value, provide an explanation as to why you labelled this predicate as having this particular value. Use the format: <<predicate>>: <<truth_value>>. <<explanation>>. Predicates:
We then specify the ground atom strings whose values we’d like to query. For instance, in one of the Burger tasks, some of these might be:
@@
Two aspects of our prompting strategy proved essential for achieving high labeling accuracy. First, chain-of-thought prompting [49] was extremely useful: our prompt (for $t>0$) asks the VLM to explicitly provide reasoning for its choices in addition to labeling the truth value. Prompting the VLM to describe differences between two scenes in a chain-of-thought fashion led to much more accurate output than simply labeling atoms based on a single scene. Second, we found that a form of set-of-marks prompting [50] where we annotate objects in the image with their names to be critical to the VLM being able to correctly identify objects being referred to in the text prompt. We only perform this annotation in the simulation environments, as we found it to be unnecessary and additional engineering effort in the real-world environments. Thus, in addition to assuming we have been given a set of user-specified objects (with names, an optional type, and descriptor string), we also assume, for the simulation environments (in which all objects have a type), that we have access to an object detection system that can classify objects according to their type, and track individual object instances over time, disambiguating between multiple objects of the same type across frames. An example of such annotations in the Burger and Coffee environments is shown in Figure 5. Notice that we differentiate between ”cup0” and ”cup1”. In the simulation environments, an annotation for an object is just the type of the object plus a number.
Figure 5: Example object annotations in Burger and Coffee.
@@
For all states $s_{t}$ in a trajectory that are not the initial state (i.e., $t>0$), we prompt the VLM with the following information: (1) the image-based state from the current and previous timesteps (i.e. $s_{t}^{\text{img}}$ and $s_{t-1}^{\text{img}}$), (2) the action executed to get to the current state (i.e. $a_{t-1})$, and (3) the response of the VLM from the previous timestep. Here is an example of a prompt in one of the Burger tasks, at particular timestep.
@@
You are a vision system for a robot. You are provided with two images corresponding to the states before and after a particular skill is executed. You are given a list of predicates below, and you are given the values of these predicates in the image before the skill is executed. Your job is to output the values of the following predicates in the image after the skill is executed. Pay careful attention to the visual changes between the two images to figure out which predicates change and which predicates do not change. For the predicates that change, list these separately at the end of your response. Note that in some scenes, there might be no changes. First, output a description of what changes you expect to happen based on the skill that was just run, explicitly noting the skill that was run. Second, output a description of what visual changes you see happen between the before and after images, looking specifically at the objects involved in the skill’s arguments, noting what objects these are. Next, output each predicate value in the after image as a bulleted list with each predicate and value on a different line. For each predicate value, provide an explanation as to why you labeled this predicate as having this particular value. Use the format: <<predicate>>: <<truth_value>>. <<explanation>>. Your response should have three sections. Here is an outline of what your response should look like: [START OUTLINE] # Expected changes based on the executed skill [insert your analysis on the expected changes you will see based on the skill that was executed] # Visual changes observed between the images [insert your analysis on the visual changes observed between the images] # Predicate values in the after image [insert your bulleted list of ‘* ¡predicate¿: ¡truth value¿. ¡explanation¿‘] [END OUTLINE] Predicates: available(robot) busy(robot) cooked(patty1) empty(bottom_bun1) empty(grill) empty(robot) empty_grill(grill) free(robot) Skill executed between states: Place[robot:robot, top_bun1:top_bun, patty1:patty] Predicate values in the first scene, before the skill was executed: # Expected changes based on the executed skill The skill executed is ”Pick[robot:robot, top_bun1:top_bun]”. We expect the robot to pick up the top bun. The top bun should no longer be on the ground or table and should be in the robot’s possession. # Visual changes observed between the images In the before image, the top bun is on the ground. In the after image, the robot is holding the top bun, indicating that the robot has successfully picked it up. # Predicate values in the after image * available(robot): Unknown. The previous value was unknown, and the pick action does not provide information about availability. * busy(robot): Unknown. The previous value was unknown, and the pick action does not provide information about busyness. * cooked(patty1): True. The patty remains cooked, as it was cooked before the action. Previously, the value was true. * empty(bottom_bun1): False. The bottom bun remains unchanged. Previously, the value was false. * empty(grill): True. The grill remains empty, as the action does not involve the grill. Previously, the value was true. * empty(robot): False. The robot is now holding the top bun. Previously, the value was true. * empty_grill(grill): True. The grill remains empty, as the action does not involve the grill. Previously, the value was true. * free(robot): Unknown. The previous value was unknown, and the pick action does not provide information about freedom. …
@@
There are two additional prompting techniques we found helpful and important for labeling accuracy: (1) asking the VLM to “double-check” its label output based on the explanations it provided, and (2) augmenting the image-based state to include a close-up crop of the image for the objects the robot is currently interacting with. We perform (2) only in the Burger domain. For double-checking, we provide the VLM with the following prompt:
@@
We use double-checking only at training time to maximize labeling accuracy before optimization.
A.5.2 Atom Proposal
@@
Predicates for Each Action 1. Pick[robot:robot, patty1:patty] Before: on(patty1, cutting_board) After: holding(robot, patty1) Synonyms: grasping(robot, patty1), carrying(robot, patty1) Antonyms: releasing(robot, patty1), dropping(robot, patty1) 2. Place[robot:robot, patty1:patty, grill:grill] Before: holding(robot, patty1) After: on(patty1, grill) Synonyms: atop(patty1, grill), above(patty1, grill) Antonyms: off(patty1, grill), under(patty1, grill) 3. Cook[robot:robot, patty1:patty, grill:grill] Before: raw(patty1) After: cooked(patty1) Synonyms: grilled(patty1), prepared(patty1) Antonyms: raw(patty1), uncooked(patty1) 4. Pick[robot:robot, patty1:patty] Before: on(patty1, grill) After: holding(robot, patty1) Synonyms: grasping(robot, patty1), carrying(robot, patty1) Antonyms: releasing(robot, patty1), dropping(robot, patty1) 5. Place[robot:robot, patty1:patty, bottom_bun1:bottom_bun] Before: holding(robot, patty1) After: on(patty1, bottom_bun1) Synonyms: atop(patty1, bottom_bun1), above(patty1, bottom_bun1) Antonyms: off(patty1, bottom_bun1), under(patty1, bottom_bun1) 6. Pick[robot:robot, top_bun1:top_bun] Before: on(top_bun1, cutting_board) After: holding(robot, top_bun1) Synonyms: grasping(robot, top_bun1), carrying(robot, top_bun1) Antonyms: releasing(robot, top_bun1), dropping(robot, top_bun1) 7. Place[robot:robot, top_bun1:top_bun, patty1:patty] Before: holding(robot, top_bun1) After: on(top_bun1, patty1) Synonyms: atop(top_bun1, patty1), above(top_bun1, patty1) Antonyms: off(top_bun1, patty1), under(top_bun1, patty1) Other Important Predicates Additional Initial State Predicates: on(bottom_bun1, cutting_board) on(cheese1, cutting_board) on(lettuce1, cutting_board) near(robot, grill) near(robot, bottom_bun1) near(robot, top_bun1) Synonyms: on(x, y): atop(x, y), above(x, y) holding(x, y): grasping(x, y), carrying(x, y) cooked(x): grilled(x), prepared(x) raw(x): uncooked(x), fresh(x) near(x, y): close_to(x, y), adjacent_to(x, y) Antonyms: on(x, y): off(x, y), under(x, y) holding(x, y): releasing(x, y), dropping(x, y) cooked(x): raw(x), uncooked(x) raw(x): cooked(x), prepared(x) near(x, y): far_from(x, y), distant_from(x, y)
We do the same set-of-marks prompting that we do for atom labeling, for atom proposal. Also, note that in the Coffee environment, we add ”Generate only 5-10 predicates” to the proposal prompt to help restrict the total number of predicates the VLM proposes. This alleviates some of the difficulty the hill-climbing procedure would normally have to deal with, when faced with a very large set of predicates (e.g. 200+). Note that the VLM proposes atoms for each trajectory, so when we ask it to generate 5-10 atoms, it generates between $5n$ and $10n$ atoms in total if we had n demonstrations.
@@ A.6 Additional Experimental Details
A.6.1 Experimental Evaluation and Hyperparameters.
We used the following hyperparameter values in our different domains. Where applicable, these values were determined via a validation set. Here $N_{\text{demo}}$ is the size of the training dataset.
Task $J_{\text{thresh}}$ $h_{\text{pre\_frac}}$ $h_{\text{data\_frac}}$ $N_{\text{demo}}$
Kitchen 100 0.8 0.05 3
Bigger Burger 2000 0.8 0.05 12
More Burger Stacks 2000 0.8 0.05 12
Combo Burger 2000 0.8 0.05 12
Coffee 100 0.8 0.05 5
Cleanup 100 1.0 0.4 6
Juice 0 1.0 0.4 10
Table 2: Domain parameters used in experiments.
A.6.2 Additional Environment Details.
@@
        KettleBoiling(?k:kettle, ?kn:knob, ?b:burner): is true only if ?kn is turned on, and the kettle ?k is ontop of ?b, and if ?kn and ?b are ’linked’ (i.e., ?kn is the knob that causes burner ?b to glow red with heat).
@@
      Cook(?r:robot, ?p:patty, ?g:grill, []). Given the robot is adjacent to grill ?g and patty ?p is atop the grill, cooks the patty so it appears grilled.
@@
      Chop(?r:robot, ?l:lettuce, ?c:cutting_board, []). Given the robot is adjacent to cutting board ?c and lettuce ?l is atop the cutting board, chops the lettuce with a knife that’s on the cutting board so it appears chopped.
@@
          SomewhereAboveAndPrepped(?l:lettuce, ?b:bottom_bun): turns true when ?l is somewhere above ?b and ?l is chopped.